# The Negation Of Singer’s Conjecture For The Sixth Algebraic Transfer

### v6

@@ Abstract
Let $\mathscr{A}$ be the Steenrod algebra over the field of characteristic two, $\mathbb{F}_{2}.$ Denote by $GL(q)$ the general linear group of rank $q$ over $\mathbb{F}_{2}.$ The algebraic transfer, introduced by W. Singer [Math. Z. 202 (1989), 493-523], is a rather effective tool for unraveling the intricate structure of the (mod-2) cohomology of the Steenrod algebra, ${\rm Ext}_{\mathscr{A}}^{q,*}(\mathbb{F}_{2},\mathbb{F}_{2}).$ The Kameko homomorphism is one of the useful tools to study the dimension of the domain of the Singer transfer. Singer conjectured that the algebraic transfer is always a monomorphism, but this remains open in general case. In this paper, by constructing a novel algorithm implemented in the computer algebra system OSCAR for computing $GL(q)$-invariants of the kernel of the Kameko homomorphism, we disprove Singer’s conjecture for bidegree $(6,6+36).$
@@
1 Introduction and statement of the main outcome
@@
where $M$ is the matrix whose columns are the coordinate vectors (with respect to the monomial basis of $P_{q}$) of the degree-$n$ basis elements in $\mathscr{A}^{>0}\cdot P_{q}.$ However, obtaining a closed formula for $\operatorname{rank}(M)$—equivalently, for $\dim(QP_{q})_{n}$—for arbitrary $q$ and $n$ currently appears infeasible. It is therefore important to seek effective bounds for $\operatorname{rank}(M),$ and hence for $\dim(QP_{q})_{n}$ via the identity above. Using a new approach based on graph theory and combinatorics, our recent work [32] establishes the following result:
$LB_{match}(q,n)\leq\text{rank}(M)\leq\min\left\{\binom{n+q-1}{q-1}-\mathcal{S}_{q}(n),\ \ \mathcal{W}_{q}(n)\right\},$
where the bounding terms are explicitly computable formulas defined as follows:
  • 1.
    The ”spike” count $\mathcal{S}_{q}(n)$: This term counts the number of spike monomials. It is given by:
    (*) $\mathcal{S}_{q}(n)=\sum_{\begin{subarray}{c}(c_{0},c_{1},\dots)\in\mathbb{Z}_{\geq 0},\\ \sum c_{m}=q,\,\sum c_{m}2^{m}=q+n\end{subarray}}\frac{q!}{\prod_{m\geq 0}c_{m}!}.$
  • 2.
    The non-zero column bound $\mathcal{W}_{q}(n)$: This term provides an upper bound on the rank by summing the number of potentially non-zero columns, given by:
    $\mathcal{W}_{q}(n)=\sum_{0\leq t\leq\lfloor\log_{2}n\rfloor}\left(\binom{n-2^{t}+q-1}{q-1}-Z^{\prime}_{t}(q,n-2^{t})\right),$
    where $Z^{\prime}_{t}(q,s)$ counts monomials of degree $s$ annihilated by $Sq^{2^{t}}$.
  • 3.
    The matching-based lower bound $LB_{match}(q,n)$: This provides a lower bound for the rank based on a matching argument on the bipartite support graph of the matrix $M$, given by:
    $LB_{match}(q,n)=\left\lceil\frac{E(q,n)}{\Delta(q,n)}\right\rceil,$
    where $E(q,n)$ is the total number of non-zero entries in $M$, and $\Delta(q,n)$ is the maximum vertex degree in the support graph.
To provide readers with the context related to the inequality above, we restate the following important fact, which has been mentioned in [32, Remark 2.3]: A key element in analyzing $(QP_{q})_{n}$ is the set of ”spike” monomials—monomials where every exponent is of the form $2^{m}-1$. A classical result establishes that spike monomials do not belong to $\mathscr{A}^{>0}\cdot P_{q}$ (see also [38]). Additionally, in [18], Mothebe constructed a rather involved recursive function to enumerate all spike monomials of degree $n=2^{q-1}-q$, namely
(**) $B(q,2^{q-1}-q)$ $=q\cdot B\!\left(q-1,\,2^{q-2}-(q-1)\right)+\binom{q}{3}\,B\!\left(q-3,\,2^{q-4}-(q-3)\right)$
$\quad+\sum_{5\leq r\leq q-2}\biggl[\sum_{(b_{1},b_{2},\ldots)\in[S^{r}(q)]}\binom{q}{q-r,\,b_{1},\,b_{2},\,\ldots}\biggr]B\!\left(q-r,\,2^{q-(r+1)}-(q-r)\right)$
$\quad+\sum_{(b_{1},b_{2},\ldots)\in[S^{q-1}(q)]}\binom{q}{\,b_{1}+1,\,b_{2},\,b_{3},\,\ldots}\!$,
where $[S^{r}(q)]$ is a family generated by certain tree constructions (see [18] for details), and $B(q,2^{q-1}-q)$ denotes the number of spikes of degree $2^{q-1}-q$. However, (**1) addresses only the special degree $2^{q-1}-q$. In the formula (*1), we derive an explicit general formula that applies to arbitrary $q$ and $n$.
For comparison, our formula (*1) reproduces the values given in [18] via Mothebe’s recursion (**1):
$\mathcal{S}_{2}(0)=B(2,0)=1,\qquad\mathcal{S}_{3}(1)=B(3,1)=3,\qquad\mathcal{S}_{4}(4)=B(4,4)=13,$
$\mathcal{S}_{5}(11)=B(5,11)=75,\qquad\mathcal{S}_{6}(26)=B(6,26)=525,\qquad\mathcal{S}_{7}(57)=B(7,57)=4,347.$
Remarkably, Mothebe [18] gives a worked example for $q=11$ and $n=1013$ with the claim $B(11,1013)=135,029,697$. Nevertheless, this hand computation is inaccurate. In [28], we implemented a SageMath algorithm based on Mothebe’s method and obtained $B(11,1013)=68,958,747$. Applying our closed form (*1) independently yields the same value, $\mathcal{S}_{11}(1013)=68,958,747$. This shows that the hand calculation in [18] for $q=11$, $n=1013$ is not true.
@@
Over the past nearly four decades, the Singer transfer and related aspects have been extensively studied by numerous authors (see, e.g., [1, 3, 7, 8, 9, 10, 11, 12, 13, 17, 19, 21, 22, 34, 35, 36, 24, 25, 26, 27, 28, 29, 30, 31, 32, 37, 38, 39, 40]). In particular, in [1], Boardman showed that $Tr_{3}(\mathbb{F}_{2})$ is also an isomorphism. Remarkably, Singer [33] proved that the algebraic transfer fails to be surjective in bidegree $(5,14),$ after which he proposed the following ensuing unsettled conjecture.
@@
The conjecture is also very difficult to attack, partly because the calculation of both the domain and the codomain of $Tr_{q}(\mathbb{F}_{2})$ is not easy. It has remained an open problem for over three decades when $q\geq 4.$ Our recent works, as presented in [24, 25, 26], have successfully confirmed the conjecture’s validity for $q=4.$ In this paper, we show that the conjecture does not remain valid for the $q=6$ case.
Statement of the main outcome. We refute Conjecture 1.1 for bidegree $(6,6+36)$ by explicitly determining both the dimensions of the domain and codomain of $Tr_{6}(\mathbb{F}_{2}).$ (Note that $\mu(36)=4<6.$) We obtain the following.
@@
For instance, let us consider the case $q=5,\,n=35.$ In [35], Nguyen Sum had only determined the dimension of the invariant space $[(QP_{5})_{35}]^{GL(5)},$ to be one, without providing an explicit basis. Our new algorithm’s output not only confirms this dimension but also furnishes an explicit basis for this space (including the dimension and basis for the invariant space of the kernel of the Kameko homomorphism). In particular, for $q=5$ and degree $35,$ our algorithm finds that the invariant space of the kernel of the Kameko homomorphism is trivial, while the invariant space $[(QP_{5})_{35}]^{GL(5)}$ is one-dimensional. The algorithm further shows that $[(QP_{5})_{35}]^{GL(5)}=\mathbb{F}_{2}\cdot[\texttt{GL5[1]}],$ where
@@
Here $\psi$ is the homomorphism $\psi:(P_{5})_{15}\longrightarrow(P_{5})_{35},\,x_{1}^{e_{1}}\ldots x_{5}^{e_{5}}\longmapsto x_{1}^{2e_{1}+1}\ldots x_{5}^{2e_{5}+1},$ and the polynomial $q$ is determined as in Subsection 6.6 of [35]. Re-verifying the above result by hand is also not too difficult. For the reader’s convenience, we also provide detailed output of our algorithm for the case $q=5,\,n=35$ at:
https://drive.google.com/file/d/1qyQOV2RX23afcWhwzNdLfFBHF-5SiUCm/.
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [36], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [22, 23]). As a limitation of this approach, the result in [36] remains unverified by modern computer algebra systems such as OSCAR, Magma, or SageMath. In fact, verifying all the manual computations (now considered outdated) in [36] on a computer algebra system is very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [30]). This implies that the manual computational methods in [36], while not novel, become impractical at higher ranks, which explains why Singer’s Conjecture 1.1 remained open until this work. As discussed above, we address this limitation for the $(q,n)=(6,36)$ case in Theorem 1.4, where we provide full computational verification via computer algebra systems.
We also want to emphasize a key point about computations for the space $(QP_{q})_{n}$ and its invariants. While manual calculations can be verified in some low-degree cases, verification for higher degrees, such as $q=5,\,n=108,$ is only feasible on a computer algebra system. The reason is that the number of input monomials, determined by the formula $\binom{n+q-1}{q-1}$, grows enormously with the degree $n,$ making the task of manually checking results practically impossible.
@@ 2 A few preliminaries
As discussed in Section 1, our focus is on understanding both the behavior of the Singer algebraic transfer and Conjecture 1.1. In particular, the Singer conjecture is essential for studying the structure of the cohomology groups of the Steenrod algebra. To address this conjecture for the $q=6$ case, we need to explicitly determine the domain and codomain of the transfer map $Tr_{6}(\mathbb{F}_{2}).$ Remarkably, the domain of $Tr_{6}(\mathbb{F}_{2})$ is closely related to the problem of explicitly determining the dimension of the space $QP_{6}$ in positive degree $n.$ This issue is essentially about describing a minimal set of generators for the $\mathscr{A}$-module $P_{6}$, which is commonly referred to as the Peterson hit problem [21]. (For more perspectives on this remarkably difficult hit problem, we refer readers, for example, to our latest works [27, 28].) Furthermore, it is well-known that the domain of $Tr_{q}(\mathbb{F}_{2})$ is dual to the $GL(q)$-invariant $[(QP_{q})_{n}]^{GL(q)}$ for any positive degree $n.$ Therefore, determining $\mathscr{A}$-generators for $P_{q}$ at degree $n$ stands as a crucial undertaking. Building on this relationship, we need to consider the following concepts.
@@
3 Proof of Theorem 1.4 using algorithms in SageMath and OSCAR
@@
In [28], we showed that $\dim(QP_{6})_{15}=2184.$ So, we need only to determine $\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ to deduce the dimensional result for $(QP_{6})_{36}.$ And from that, based on a basis for $(QP_{6})_{36}$ and the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 6,$ we can explicitly compute the dimension and basis for the invariant space $[(QP_{6})_{36}]^{GL(6)}.$
We notice that computing by hand an explicit basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ is a hard and error-prone task, due to the growing number of monomials as the number of variables and degrees increases. Manual computation can typically be controlled in cases where the degree is not too large and the increase in the number of monomials is manageable (for instance, one can see some works by the author [26, 27, 28], the author and Nguyen Sum [22], and Nguyen Sum [34] to understand the specific manual computation methods). Therefore, in recent works [29, 30, 31], we have developed computational programs to explicitly compute the dimension of $(QP_{q})_{n}$ and the invariant $[(QP_{q})_{n}]^{GL(q)}.$ These algorithmic programs allow us to display detailed computations as readers have become familiar with in our previous works, and some other authors.
@@
    https://drive.google.com/file/d/190UNigq7PtKasrcu3qg44_2Sqr_qqH0P/
@@
    https://drive.google.com/file/d/14n4wXo0lYP8ciPMyMmrBH2CGiCyW1dcZ/

### v7

@@ Abstract
Let $\mathscr{A}$ be the Steenrod algebra over the field of characteristic two, $\mathbb{F}_{2}.$ Denote by $GL(q)$ the general linear group of rank $q$ over $\mathbb{F}_{2}.$ The algebraic transfer, introduced by W. Singer [Math. Z. 202 (1989), 493-523], is a rather effective tool for unraveling the intricate structure of the (mod-2) cohomology of the Steenrod algebra, ${\rm Ext}_{\mathscr{A}}^{q,*}(\mathbb{F}_{2},\mathbb{F}_{2}).$ The Kameko homomorphism is one of the useful tools to study the dimension of the domain of the Singer transfer. Singer conjectured that the algebraic transfer is always a monomorphism. This conjecture has stood for nearly 40 years, and in this work we demonstrate that it fails in general. More precisely, by developing a novel algorithm in the computer algebra system OSCAR for computing $GL(q)$-invariants of the kernel of the Kameko homomorphism, we disprove Singer’s conjecture for bidegree $(6,6+36)$.
@@
1 Introduction and statement of the main result
@@
where $M$ is the matrix whose columns are the coordinate vectors (relative to the monomial basis of $P_{q}$) of the degree-$n$ basis elements in $\mathscr{A}^{>0}\cdot P_{q}$. However, finding a closed formula for $\operatorname{rank}(M)$, and hence for $\dim(QP_{q})_{n}$, for general $q$ and $n$ appears to be infeasible at present. It is therefore crucial to find effective bounds on $\operatorname{rank}(M)$, which would yield bounds on $\dim(QP_{q})_{n}$ through the identity above. We refer the reader to our recent article [30], where this problem is treated and an explicit bound for $\dim(QP_{q})_{n}$ is derived.
@@
Over the past nearly four decades, the Singer transfer and related aspects have been extensively studied by numerous authors (see, e.g., [1, 3, 7, 8, 9, 10, 11, 12, 13, 17, 18, 20, 21, 32, 33, 23, 26, 27, 28, 29, 30, 36, 37, 38]). In particular, in [1], Boardman showed that $Tr_{3}(\mathbb{F}_{2})$ is also an isomorphism. Remarkably, Singer [31] proved that the algebraic transfer fails to be surjective in bidegree $(5,14),$ after which he proposed the following ensuing unsettled conjecture.
@@
The conjecture is also very difficult to attack, partly because the calculation of both the domain and the codomain of $Tr_{q}(\mathbb{F}_{2})$ is not easy. The conjecture is known to hold for $q=1,\,2$ by Singer [31], for $q=3$ by Boardman [1], and for $q=4$ by our recent works [23, 24, 25]. In this paper, we demonstrate that it does not hold in general.
Statement of our main result. We refute Conjecture 1.1 in general by showing it fails in bidegree $(6,6+36)$. (Note that $\mu(36)=4<6$.) We proceed by explicitly computing the dimensions of the domain and codomain of $Tr_{6}(\mathbb{F}_{2})$ in internal degree $36$. The following results are obtained.
@@
For instance, let us consider the case $q=5,\,n=35.$ In [33], Nguyen Sum had only determined the dimension of the invariant space $[(QP_{5})_{35}]^{GL(5)},$ to be one, without providing an explicit basis. This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [21, 22]). Such manual computations have since become obsolete. Our algorithm not only confirms this dimension but also provides an explicit basis for this space, including the dimension and basis of the invariant subspace of the kernel of the Kameko homomorphism. In particular, for $q=5$ and degree $35,$ our algorithm finds that the invariant space of the kernel of the Kameko homomorphism is trivial, while the invariant space $[(QP_{5})_{35}]^{GL(5)}$ is one-dimensional. The algorithm further shows that $[(QP_{5})_{35}]^{GL(5)}=\mathbb{F}_{2}\cdot[\texttt{GL5[1]}],$ where
@@
Here $\psi$ is the homomorphism $\psi:(P_{5})_{15}\longrightarrow(P_{5})_{35},\,x_{1}^{e_{1}}\ldots x_{5}^{e_{5}}\longmapsto x_{1}^{2e_{1}+1}\ldots x_{5}^{2e_{5}+1},$ and the polynomial $q$ is determined as in Subsection 6.6 of [33, arXiv:1609.03006]. Re-verifying the above result by hand is also not too difficult. For the reader’s convenience, we also provide detailed output of our algorithm for the case $q=5,\,n=35$ at: https://doi.org/10.5281/zenodo.17620639.
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [34], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [21, 22]). As a limitation of this approach, the result in [34] remains unverified by modern computer algebra systems such as OSCAR, Magma, or SageMath. In fact, verifying all the manual computations (now considered outdated) in [34] on a computer algebra system is very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [28]). This implies that the manual computational methods in [34], while not novel, become impractical at higher ranks, which explains why Singer’s Conjecture 1.1 remained open until our present work. As discussed above, we address this limitation for the $(q,n)=(6,36)$ case in Theorem 1.4, where we provide full computational verification via computer algebra systems.
@@ 2 A few preliminaries
As discussed in Section 1, our focus is on understanding both the behavior of the Singer algebraic transfer and Conjecture 1.1. In particular, the Singer conjecture is essential for studying the structure of the cohomology groups of the Steenrod algebra. To address this conjecture for the $q=6$ case, we need to explicitly determine the domain and codomain of the transfer map $Tr_{6}(\mathbb{F}_{2}).$ Remarkably, the domain of $Tr_{6}(\mathbb{F}_{2})$ is closely related to the problem of explicitly determining the dimension of the space $QP_{6}$ in positive degree $n.$ This issue is essentially about describing a minimal set of generators for the $\mathscr{A}$-module $P_{6}$, which is commonly referred to as the Peterson hit problem [20]. (For more perspectives on this remarkably difficult hit problem, we refer readers, for example, to our latest works [26].) Furthermore, it is well-known that the domain of $Tr_{q}(\mathbb{F}_{2})$ is dual to the $GL(q)$-invariant $[(QP_{q})_{n}]^{GL(q)}$ for any positive degree $n.$ Therefore, determining $\mathscr{A}$-generators for $P_{q}$ at degree $n$ stands as a crucial undertaking. Building on this relationship, we need to consider the following concepts.
@@
3 Proof of Theorem 1.4
@@
In [Phuc4], we showed that $\dim(QP_{6})_{15}=2184.$ So, we need only to determine $\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ to deduce the dimensional result for $(QP_{6})_{36}.$ And from that, based on a basis for $(QP_{6})_{36}$ and the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 6,$ we can explicitly compute the dimension and basis for the invariant space $[(QP_{6})_{36}]^{GL(6)}.$
We notice that computing by hand an explicit basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ is a hard and error-prone task, due to the growing number of monomials as the number of variables and degrees increases. Manual computation can typically be controlled in cases where the degree is not too large and the increase in the number of monomials is manageable (for instance, one can see some works by the author [25, 26], the author and Nguyen Sum [21], and Nguyen Sum [32] to understand the specific manual computation methods). Therefore, in recent works [27, 28, 29], we have developed computational programs to explicitly compute the dimension of $(QP_{q})_{n}$ and the invariant $[(QP_{q})_{n}]^{GL(q)}.$ These algorithmic programs allow us to display detailed computations as readers have become familiar with in our previous works, and some other authors.
@@
    https://doi.org/10.5281/zenodo.17620661.
@@
    https://doi.org/10.5281/zenodo.17620680.
@@
Declaration of Competing Interests
The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.